OpenStudy (marysammonss):

Which equation has solution x = -3? A) 2x - 7 = -1 Eliminate B) 3x + 8 = 1 C) 1 2 x + 8 = 10 D) 1 2 (2x - 6) = -6

OpenStudy (jackthegreatest):

plug in -3 into the equation

OpenStudy (jackthegreatest):

for example: 2 (-3) - 7 = -1

OpenStudy (jackthegreatest):

Is it possible to choose none of the above?

OpenStudy (marysammonss):

No, i really dont know maybe its just the test lol

OpenStudy (jackthegreatest):

make sure you listed the equations right, because none of them have solutions of -3...

OpenStudy (marysammonss):

Which equation has solution x = -3? A) 2x - 7 = -1 B) 3x + 8 = 1 C) 1/2x + 8 = 10 D) 1/2(2x - 6) = -6

OpenStudy (marysammonss):

i think i typed it wrong

OpenStudy (jackthegreatest):

ahh yes that makes a difference

OpenStudy (jackthegreatest):

the answer is D because: 1/2 (2(-3) -6) = -6 1/2 (-6-6) = -6 1/2 (-12) = -6 -6=-6

OpenStudy (jackthegreatest):

@marysammonss